Even though most corporate bonds in the United States make coupon payments semiannually, bonds issued elsewhere often have annual coupon payments. Suppose a German company issues a bond with a par value of €1,000, 15 years to maturity, and a coupon rate of 6.5 percent paid annually. If the yield to maturity is 7.6 percent, what is the current price of the bond? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Coupon = 0.065 * 1000 = 65

Current price = Coupon* [1 - 1 / (1 + r)n] / r + FV / (1 +r)n

Current price = 65* [1 - 1 / (1 + 0.076)15] / 0.076 + 1000 / (1 +0.076)15

Current price = 65* [1 - 0.333285] / 0.076 + 333.285124

Current price = 65* 8.772566 + 333.285124

Current price = $903.50